### v1

# Large Language Models in Argument Mining: A Survey

Abstract
Argument Mining (AM), a critical subfield of Natural Language Processing (NLP), focuses on extracting argumentative structures from text. The advent of Large Language Models (LLMs) has profoundly transformed AM, enabling advanced in-context learning, prompt-based generation, and robust cross-domain adaptability. This survey systematically synthesizes recent advancements in LLM-driven AM. We provide a concise review of foundational theories and annotation frameworks, alongside a meticulously curated catalog of datasets. A key contribution is our comprehensive taxonomy of AM subtasks, elucidating how contemporary LLM techniques—such as prompting, chain-of-thought reasoning, and retrieval augmentation—have reconfigured their execution. We further detail current LLM architectures and methodologies, critically assess evaluation practices, and delineate pivotal challenges including long-context reasoning, interpretability, and annotation bottlenecks. Conclusively, we highlight emerging trends and propose a forward-looking research agenda for LLM-based computational argumentation, aiming to strategically guide researchers in this rapidly evolving domain.
This work is in progress. The content and references are subject to change.
1 Introduction
Argument Mining (AM) is a fast-growing subfield of Natural Language Processing (NLP) concerned with the automatic identification and extraction of argumentative structures—claims, premises, and the relations between them—from textual discourse (Lawrence & Reed, 2019; Patel, 2024). Such structure is central to understanding not just what opinions are expressed but why, enabling applications ranging from debate moderation and fact checking to legal analytics and scientific discourse analysis. Indeed, by transforming unstructured text into structured argument representations, AM supports reasoning-centered applications that go far beyond traditional sentiment analysis, which can determine opinions but not their underlying rationale (Lawrence & Reed, 2019).
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio & Villata, 2018). Foundational surveys—most notably those by Lawrence & Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence & Reed, 2019); Peldszus & Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus & Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
Building on these foundations, subsequent studies have enriched the field with domain- and method-specific insights. Cabrio & Villata (2018) offered a data-driven reflection on the field’s early corpus and model evolution , while targeted surveys explored annotation practices, tool ecosystems, and argument quality assessment. For example, Patel (2024) traces the integration of traditional machine learning with emerging neural techniques; Schaefer & Stede (2021) focus on Twitter; Hua et al. (2019) and Fromm et al. (2021) examine peer-review text; and Guida et al. (2025) assess the performance of LLMs on enforcing topic-specific argument tasks in online comments (Altemeyer et al., 2025; Chen et al., 2024).
However, despite this breadth, most existing reviews remain anchored in the pre‑LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, and reconnects these technical developments to foundational argumentation theory and annotation practice.
In response, this survey offers a timely and theory-informed synthesis of how modern Large Language Models are redefining argument mining. We present AM not as a linear sequence of isolated subtasks but as a network of interdependent LLM-powered components, united by prompt engineering, chain-of-thought reasoning, few-shot learning, and retrieval-augmented generation. We examine these innovations across a wide range of domains—from legal and scientific text to social media, education, peer reviews, and emerging multi-modal content—surveying benchmark results from the past three years as well as emphasizing the empirical gaps in long-context reasoning, multilingual adaptation, model interpretability, and efficient deployment.
Building on this foundation, the rest of the survey unfolds as follows.Section 2 catalogs datasets and benchmarks—emphasizing recent large, cross-domain, and multimodal resources suited for LLM evaluation. Section 3 develops a taxonomy of key AM subtasks—claim detection, premise identification, relation parsing, quality assessment, and summarization—exploring how LLM techniques like prompting, chain‑of‑thought, few‑shot learning, and retrieval‑augmented generation reshape each. Section 4 introduces model architectures and key LLM-era techniques, detailing encoder–decoder designs, prompting strategies, fine-tuning approaches, and multimodal integration relevant to argument mining. Section 5 outlines evaluation methodologies, including metrics, human-in-the-loop assessments, and benchmarking practices tailored to AM. Section 6 delves into ongoing challenges—such as long-context reasoning, domain transfer, interpretability, annotation bottlenecks, and computational efficiency—while highlighting emerging research trends. Finally, Section 7 presents our conclusions and proposes a forward-looking roadmap for advancing LLM-powered computational argumentation.
2 Dataset and Benchmark
The landscape of resources for argument mining has undergone a significant evolution, marked by a clear trajectory from the analysis of static, decontextualized text towards a more nuanced understanding of argumentation in its natural settings. This progression is evidenced by the creation of novel datasets, particularly between 2021 and 2024, which not only expand the volume of available data but also fundamentally advance the complexity of tasks and the scope of inquiry within the field. These resources increasingly prioritize the dynamic, interactive, and multimodal facets of argumentation, reflecting the field’s growing maturity.
2.1 Capturing Argumentation in Dialogue
A primary vector of recent dataset development has been the focus on conversational argumentation. Recognizing that real-world arguments are predominantly forged through interaction, recent efforts have sought to capture the intricate dynamics of dialogue. A significant development in this area is the ADEA (Argumentative Dialogue Dataset on Ethical Issues Concerning Future A.I. Applications) Hauptmann et al. (2024). This dataset provides dialogues centered on the complex and value-laden topic of AI ethics, enabling research into how arguments are proposed, contested, and defended within a specialized, high-stakes domain.
Moving towards a more granular analysis of dialogue mechanics, the CAMAL (Conversational Argument Move Analysis) dataset introduces a multi-label framework for annotating ”conversational moves” Lai et al. (2024). This approach moves beyond the traditional support/attack dichotomy to capture the multifaceted functions of utterances, such as questioning, clarifying, or challenging a premise. Such a fine-grained annotation scheme more accurately reflects the fluidity of natural dialogue and supports the development of models for more sophisticated dialogical systems. The challenge of parsing multi-party dialogue is addressed by the work of Mao et al., who developed a new argumentative dataset alongside a hierarchical graph fusion network Mao et al. (2024). Their contribution underscores the necessity of understanding global discourse structure for accurately resolving local argumentative relations in complex, multi-participant settings. Reinforcing this theme, the work of Mezza et al. demonstrates that exploiting dialogue acts and contextual information significantly improves the identification of argumentative relations in online debates Mezza et al. (2024).
2.2 Multi-modality Benchmark
As the number of specialized datasets has grown, the issue of fragmentation has become more salient, prompting efforts to foster model generalization and broaden the modalities of analysis. The ARIES (A General Benchmark for Argument Relation Identification) project represents a critical step towards unification Gemechu et al. (2024). By systematically integrating multiple existing datasets from diverse domains and annotation traditions, ARIES establishes a standardized benchmark for a core argument mining task. This initiative facilitates more equitable model comparisons and encourages the development of more robust algorithms that are not over-fitted to the idiosyncrasies of a single corpus. The importance of data composition is further explored by Schiller et al., who find that for topic-dependent argument mining, dataset diversity often has a greater impact on model performance than sheer size, providing crucial guidance for future data collection efforts Schiller et al. (2024).
The field’s scope has also expanded beyond purely textual data. The M-Arg dataset was a pioneering contribution in this direction, offering a multimodal resource for argument mining in political debates that includes both audio recordings and aligned transcripts Mestre et al. (2021). This enables the joint analysis of linguistic content and acoustic-prosodic features (e.g., pitch, tone, pauses), allowing researchers to investigate how non-verbal cues contribute to the expression of argumentative functions. M-Arg thus provides the empirical foundation for building computational models capable of a more holistic understanding of argumentation.
2.3 Argument Structure
Concurrent with the expansion of scope, there is a clear trend towards deepening the analysis of core argumentative concepts and applying them in highly specialized domains. The legal field, with its inherently structured and argumentative nature, remains a key area of application. The dataset of annotated European Court of Human Rights (ECHR) cases presented by Mumford et al. offers detailed insights into the intricacies of legal reasoning Mumford et al. (2024). Similarly, in the medical domain, the CasiMedicos-Arg dataset annotates explanatory argumentative structures within medical question-answering contexts, linking argument mining directly to the critical need for explainability in clinical AI Sviridova et al. (2024a; b).
The assessment of argument quality has emerged as another major research frontier. Large-scale datasets such as ArgAnalysis35K Joshi et al. (2023) and the work by Ivanova et al. on quality dimensions Ivanova et al. (2024) signal a shift from identifying the presence of an argument to evaluating its intrinsic merit. These resources facilitate the development of models that can assess arguments along multiple axes, including logic, sufficiency, and clarity, which is essential for applications such as writing support and automated feedback systems. Furthermore, there is a renewed focus on the foundational logical and philosophical underpinnings of argumentation. The EthiX dataset for argument scheme classification in ethical debates exemplifies this, connecting computational analysis back to established argumentation theory Bezou-Vrakatseli et al. (2024). Similarly, the creation of a dataset for causality identification by Al-Khatib et al. pushes the field towards identifying the deeper reasoning patterns within argumentative texts Khatib et al. (2023).
2.4 Data Creation and Evaluation in LLM Era
The advent of Large Language Models (LLMs) has introduced a new dynamic into the creation and use of argumentative datasets. On one hand, LLMs are being leveraged as tools to overcome data scarcity. The work of de Souza and Dragoni, for instance, explores the use of LLMs to augment persuasive argument datasets, offering a scalable method for generating new training instances de Souza & Dragoni (2024). On the other hand, the argumentative capabilities of LLMs are themselves becoming a central object of study. Researchers are now creating datasets specifically designed to probe and evaluate the quality of arguments generated by models like ChatGPT. The work by Rocha et al. provides a methodology and dataset to assess ”good, bad, and ugly” machine-generated arguments Rocha et al. (2024), while Djouvas et al. present a comparative study and dataset for argument strength classification that critically evaluates the performance of LLMs against other models Djouvas et al. (2024). This reflexive turn, where LLMs are used both to create and to be evaluated by data, marks a significant new phase in the field, enabling a more nuanced understanding of the strengths and limitations of current generative models in the context of complex human reasoning.
3 Task Taxonomy
3.1 Claim Detection
Claim detection identifies central propositions requiring justification across diverse textual landscapes. This task forms the cornerstone of argument mining pipelines, with methodologies evolving to address domain-specific complexities and data modalities. The research landscape spans multiple dimensions:
3.1.1 Platform-Specific Detection Environments
Social Media & Web Platforms: These environments demand specialized approaches for informal, dynamic content. Beltrán et al. (2021) developed ClaimHunter for real-time Twitter monitoring, while Berendt et al. (2021) created FactRank for Dutch-language verification. Multimodal integration significantly enhances detection capabilities, as demonstrated by Cheema et al. (2022) through their MM-Claims dataset combining visual and textual evidence. The CLAIMSCAN framework Sundriyal et al. (2023) further advances social media truth discovery through comprehensive claim span identification.
Video & Multimodal Environments: Emerging research addresses claims in audiovisual content. Rayar (2024) pioneered multimodal video analysis techniques, while Giedemann et al. (2025) contributed ViClaim - a multilingual benchmark dataset specifically designed for video-based claim detection. These approaches must overcome unique challenges including speech-to-text conversion errors and visual-textual semantic alignment.
3.1.2 Domain-Specialized Detection Applications
Healthcare & Biomedical: This critical domain combines fraud detection Gao et al. (2019), privacy-preserving blockchain verification Mohan & Praveen (2019), and evolutionary optimization Tubishat et al. (2025). Biomedical applications focus on Twitter-based health claims Wührl & Klinger (2021) and contradictory research detection Yazi et al. (2021). Recent systematic reviews du Preez et al. (2025) evaluate machine learning efficacy across healthcare fraud patterns.
Environmental Claims: Targeting greenwashing and sustainability misinformation, researchers have developed specialized detectors Woloszyn et al. (2021) and comprehensive frameworks Stammbach et al. (2023). These systems must navigate scientific terminology and corporate reporting nuances while identifying substantiated environmental assertions.
Financial Sector: Financial claim detection addresses numerical assertions and market misinformation through systems like NumClaim Chen et al. (2020) for investor-oriented analysis. Novel datasets Shah et al. (2024) enable weak-supervised detection of numerical claims in earnings reports and market analyses, addressing domain-specific challenges like quantitative reasoning.
3.1.3 Contextual Application Scenarios
Educational Contexts: Claim detection enhances academic writing assessment and scientific discourse analysis. Wan et al. (2020) established correlations between claim presence and writing quality, while Hafid (2024) developed interpretation frameworks for scientific claims in online discourse, enabling more effective knowledge extraction from academic conversations.
Political Fact-Checking: This high-stakes domain employs innovative fairness techniques like actor masking Dayanik & Padó (2020) and positive unlabeled learning Wright & Augenstein (2020). Advanced verification frameworks Jafari & Allan (2024) improve robustness against adversarial political messaging, with rumor detection systems Lin et al. (2021) using claim-guided hierarchical attention networks to combat misinformation.
3.1.4 Cross-Lingual Methodologies
Multilingual detection addresses global misinformation through adapter fusion techniques Schlicht et al. (2023) and LLM-based approaches Vykopal et al. (2025). Cross-lingual transfer learning Panchendrarajan & Zubiaga (2025) enables knowledge sharing between languages, while transformer-based multilingual fact claiming Böck et al. (2021) demonstrates effectiveness across diverse linguistic landscapes.
3.1.5 Verification and Anomaly Detection
Claim Verification: Sophisticated validation frameworks Irnawan et al. (2025) assess claim authenticity through evidence retrieval and reasoning chains. Robust verification techniques Jafari & Allan (2024) address adversarial examples and out-of-distribution claims in high-risk domains.
Anomaly Detection: Specialized techniques identify fraudulent patterns through uncertainty modeling Vandervorst et al. (2024) and profile-based feature engineering Duval et al. (2022). These approaches analyze claim distributions across temporal, behavioral, and semantic dimensions to flag suspicious assertions in insurance, healthcare, and financial domains.
3.2 Evidence Detection
Evidence detection, a critical sub-task within argument mining and misinformation identification, involves identifying textual or multimodal components that substantiate or refute a particular claim. Recent advances in this area include diverse approaches ranging from leveraging multimodal evidence to employing graph-based methods and pretrained language models.
3.2.1 LLMs in Evidence Detection
Several studies have focused on employing pretrained language models for automated evidence extraction. Dementieva and Panchenko Dementieva & Panchenko (2021; 2020) demonstrated the effectiveness of cross-lingual evidence in bolstering monolingual fake news detection. Elaraby and Litman Elaraby & Litman (2021) employed self-trained pretrained language models specifically optimized for evidence detection, achieving notable improvements. Graph-based techniques also constitute a vital approach within evidence detection. Xu et al. Xu et al. (2022a; b) introduced an evidence-aware fake news detection method utilizing graph neural networks to mine fine-grained semantics. Moreover, Guo et al. Guo et al. (2023) developed an interpretable fake news detection model leveraging graph-based evidence, enhancing interpretability and accuracy. Interactive and dynamic evidence retrieval methodologies have also emerged. Liao et al. Liao et al. (2023a; b) proposed a multi-step evidence retrieval enhancement framework (MUSER), significantly improving accuracy by iterative evidence refinement. Furthermore, Wu et al. Wu et al. (2024) developed a unified evidence enhancement inference framework that dynamically adapts to incoming information.
3.2.2 Multilingual Evidence Detection
A significant trend involves evidence extraction across multiple sources or languages. Dementieva et al. Dementieva et al. (2023) proposed a multilingual approach named Multiverse, integrating evidence from diverse linguistic contexts. Hammouchi and Ghogho Hammouchi & Ghogho (2022) similarly developed an evidence-aware multilingual fake news detection system, underscoring the importance of linguistic diversity in robust misinformation detection.
3.2.3 Cross-domain Evidence Detection
Cross-domain adaptability and out-of-distribution scenarios have also garnered attention. Liu et al. Liu et al. (2024a) addressed out-of-distribution issues through dual adversarial debiasing, ensuring evidence-aware models remain effective when encountering novel misinformation types. Additionally, Ferdush et al. Ferdush et al. (2025) focused on cross-domain fake news detection by integrating evidence from multiple social media platforms, enhancing cross-platform robustness.
3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu & Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
3.2.5 Counter-Evidence Detection
Evidence detection is also critical for identifying contradictions and managing conflicts in factual claims. Jiayang et al. Jiayang et al. (2024) introduced the ECON framework for detecting and resolving evidence conflicts, an essential capability for maintaining informational integrity. Kharrat et al. Kharrat et al. (2022) further proposed a contradiction detection approach utilizing semantic relations and uncertainty evidence. Explainability remains integral to the credibility of evidence detection systems. Dong et al. Dong et al. (2024) proposed EMIF, an evidence-aware multi-source information fusion network emphasizing explainability, critical for transparency and trust in automated misinformation detection.
3.3 Stance Detection
Stance detection is the task of automatically identifying an author’s position towards a specific target, claim, or topic — typically labeled as pro, con, or neutral. This task plays a critical role in argument mining, especially in opinionated domains such as social media, online debates, and political discourse.
Datasets and Evaluation.
Recent efforts have produced a variety of stance detection datasets to support research in this area. For instance, Niu et al. (2025a; b) introduce C-MTCSD, a Chinese multi-turn conversational stance detection dataset, which captures nuanced stance evolution. Gyawali et al. (2024) contribute GunStance, a dataset focusing on gun regulation debates. These datasets facilitate stance classification in diverse linguistic, political, and conversational contexts.
Classical and Neural Methods.
Early neural approaches, such as those in Hosseinia et al. (2019), leverage BiLSTM architectures to detect stance in argumentative opinions. More recent work like Wang et al. (2024) introduces a recurrent stick-breaking topic model to jointly model stance and topic clusters. Furthermore, approaches like Li et al. (2024c) simulate rumor propagation to enhance stance prediction under social influence.
Large Language Models for Stance Detection.
With the rise of LLMs, several works have evaluated or leveraged their capabilities. Zhao et al. (2024) propose ZeroStance, using ChatGPT to generate datasets and perform zero-shot stance detection. Gambini et al. (2024) conduct a systematic evaluation of various LLMs (e.g., GPT-3, BLOOM) for stance detection on social media. Complementary to this, Lee et al. (2024) introduce LOGIC, where LLM-generated rationales guide small models to improve stance reasoning.
Cross-lingual and Domain Adaptation.
Multilingual stance detection remains a challenge. Charfi et al. (2024) present a multi-dialect Arabic stance corpus and cross-domain benchmarks. Works like Alghaslan & Almutairy (2024); Shukla et al. (2024) leverage LLM fine-tuning and ensembling for Arabic stance detection tasks.
Zero-shot and Few-shot Generalization.
The promise of zero-shot stance detection has spurred further innovation. Ma et al. (2024a; b) propose Chain-of-Stance, where LLMs perform chain-of-thought style reasoning. Guo et al. (2024) incorporate structured external knowledge from LLMs to improve generalization. Zhang et al. (2024) integrate commonsense knowledge through adversarial learning.
Bias Mitigation and Calibration.
Acknowledging LLM bias, Li et al. (2025; 2024a) propose counterfactual augmented calibration strategies to reduce stance-related biases in LLM outputs.
Social Media and Health Applications.
Beyond generic domains, stance detection has been used in medical and health misinformation contexts. Davydova et al. (2024) analyze COVID-19 tweet stances; Sun et al. (2024) study HPV vaccine stance detection using in-context learning and fine-tuning. Akash et al. (2024) assess the feasibility of open-target stance detection using LLMs.
Synthetic and LLM-generated Data.
LLM-generated synthetic datasets are also explored. For example, Wagner et al. (2025) show that LLM-synthesized stance data can outperform traditional fine-tuning when training low-resource classifiers. Similarly, Fraile-Hernandez & Peñas (2024) demonstrate the use of generative LLMs in political activism scenarios.
Language-independent Embeddings.
Finally, de Landa & Agerri (2025) present language-agnostic methods using interaction-based embeddings and multilingual LLMs to promote cross-lingual stance transfer.
In sum, the task of stance detection has evolved from simple text classification to a rich and multi-faceted research area, with contributions spanning LLM prompting, data generation, bias handling, and multilingual adaptability.
3.4 Argument Summarization
Argument summarization focuses on generating concise and informative summaries of argumentative texts, including debates, discussions, legal decisions, and online forums. Compared to general summarization, this task presents unique challenges due to the complexity of argumentative structure, the presence of contrasting stances, and the need to preserve key claims, evidence, and discourse relations.
Early work such as ConvoSumm(Fabbri et al., 2021) laid the foundation for integrating argument mining with abstractive summarization in dialogues, demonstrating the benefits of identifying argumentative components in conversation summarization. Building on this, several works have extended the scope of argument summarization to new domains such as law(Elaraby et al., 2023; Elaraby & Litman, 2022; Xu & Ashley, 2023; Steffes & Rataj, 2022), where legal documents are often long and require nuanced treatment of argumentative structure. The ArgLegalSumm model(Elaraby & Litman, 2022) proposed incorporating argument mining to enhance legal document summaries, and was further extended with reranking modules to improve informativeness and coherence(Elaraby et al., 2023).
In terms of modeling techniques, Shan and Lu(Shan & Lu, 2025) applied multi-task learning with argument relation-aware contrastive learning to better capture opinion divergence in debates. Similarly, Syed et al.(Syed et al., 2023) introduced frame-oriented summarization to capture pragmatic aspects of arguments in discussions. Other efforts, such as by Meer et al.(van der Meer et al., 2024), analyzed the diversity of generated summaries, revealing trade-offs between informativeness and representation of varied perspectives.
Dataset construction remains critical for progress. OpenDebateEvidence(Roush et al., 2024) provides a massive-scale benchmark with paired argument and evidence annotations, enabling more robust training. Likewise, ORCHID(Zhao et al., 2023) and ConvoSumm(Fabbri et al., 2021) support summarization across argumentative dialogues in Chinese and English, respectively.
3.4.1 Key Point Analysis
Key Point Analysis (KPA) is a fine-grained subtask within argument summarization that focuses on identifying representative and recurring ideas—referred to as “key points”—from a collection of arguments. Introduced by Bar-Haim et al. (Bar-Haim et al., 2020), KPA provides a structured framework for summarizing public opinions and argumentative discourse in a more interpretable and scalable manner than free-form summarization.
Initial approaches relied on clustering or semantic similarity. For example, Bar-Haim et al.(Bar-Haim et al., 2020; 2021) introduced the idea of extracting key points and matching arguments to them using supervised and unsupervised approaches, which inspired multiple shared tasks. Phan et al.(Phan et al., 2021) proposed a simple yet highly accurate model using bi-encoders for key point matching, while Reimer et al. (Reimer et al., 2021) explored pretrained encoders and alignment strategies.
Later efforts integrated contrastive learning (Alshomary et al., 2021) and graph-based representations (Li et al., 2024d) to improve robustness and scalability. Incremental and real-world-ready systems were also developed, as seen in Eden et al. (Eden et al., 2023).
With the rise of LLMs, researchers explored their generative and representational capabilities. Tang et al.(Tang et al., 2024b; a) examined aspect-based and prompted KPA for review summarization, highlighting the flexibility of LLMs in domain adaptation. Li et al.(Li et al., 2023) proposed an iterative clustering and abstractive summarization framework to automatically induce key points. More recently, Fukuma et al. (Fukuma et al., 2024) demonstrated that zero-shot KPA is feasible through LLM-based question generation and network analysis.
Overall, KPA bridges summarization, clustering, and representation learning. It has proven effective in condensing opinion-rich texts and continues to be enriched by LLM capabilities and multi-step reasoning frameworks.
4 Modelling Paradigms in the LLM Era
With the advent of Large Language Models (LLMs), the landscape of argument mining has undergone a significant shift. Traditional pipeline-based or shallow neural architectures are increasingly being complemented—or replaced—by foundation models that exhibit strong generalization, reasoning, and few-shot learning capabilities. This chapter categorizes the emerging modelling paradigms into four key directions: prompt-based adaptation, fine-tuning and instruction-tuning, retrieval-augmented generation, and synthetic supervision.
4.1 Prompt-based Argument Mining
Prompting has become a dominant paradigm in adapting LLMs for argument mining, enabling zero-shot and few-shot performance without task-specific supervision. Many recent works demonstrate the effectiveness of simple or chain-of-thought prompts for various sub-tasks. For example, in stance detection, Zhao et al. (2024) introduced ZeroStance, a framework that uses ChatGPT to both generate data and perform inference in a zero-shot setting. Similarly, Ma et al. (2024b) proposed Chain-of-Stance, a prompting scheme to induce intermediate reasoning steps for improved stance inference.
Prompting has also been used to assess the sufficiency or quality of arguments. Li et al. (2025) explore prompt-based counterfactual calibration strategies to mitigate stance-related biases in LLMs. Chen & Eger (2025) use prompts for manipulation checks in evaluating the effect of emotional content on argument convincingness.
These studies underscore prompting’s flexibility, interpretability, and rapid adaptability across domains and tasks.
4.2 Instruction Tuning and Multi-task Fine-tuning
While prompting alone can offer promising zero-shot performance, several works have proposed fine-tuning LLMs using argument-specific instructions or multi-task objectives. Li et al. (2024b) introduced a unified framework trained on multi-task objectives covering key point generation, evidence detection, and argument quality estimation. Their model demonstrated strong generalization across argument summarization and evaluation benchmarks.
In the legal domain, instruction-tuned models such as ArgLegalSumm (Elaraby & Litman, 2022) improve summarization by aligning model outputs with argumentative structures. Elaraby et al. (2023) further enhanced legal summarization using re-ranking mechanisms driven by LLM-computed argument salience scores.
In stance detection, models such as LOGIC (Lee et al., 2024) demonstrate that LLM-generated rationales can guide small student models through intermediate supervision—blending instruction tuning with model distillation.
4.3 Retrieval-Augmented and Knowledge-Grounded Generation
Given the inherent limitations of even the largest LLMs, retrieval-based models have emerged as a hybrid solution for argument mining. These approaches augment the model’s context window by retrieving relevant evidence, past arguments, or external knowledge.
For instance, OpenDebateEvidence (Roush et al., 2024) introduces a massive-scale retrieval-augmented dataset for claim-centric argument generation and summarization. In debate summarization, Syed et al. (2023) employ frame-oriented representations derived from retrieved background knowledge to enhance the generation of structured summaries.
Some models integrate argumentation theory into the retrieval process. Ruiz-Dolz et al. (2023) use argument graph semantics to guide the retrieval of supporting and attacking evidence, while Palmieri (2024) apply critical questions as filters during retrieval and evaluation.
4.4 Synthetic Data and LLM-generated Supervision
Synthetic supervision has emerged as a promising avenue for bootstrapping argument mining tasks at scale. LLMs can be used to synthesize high-quality training data for low-resource domains or tasks lacking annotated corpora.
Wagner et al. (2025) show that LLM-synthesized stance datasets outperform traditional few-shot fine-tuning, enabling rapid domain adaptation. Similarly, Khosravani et al. (2024) use LLM-generated key point coverage scores to construct automatic evaluation metrics for argument summarization.
This paradigm is not only cost-efficient but also enables the control and modulation of specific attributes such as diversity (van der Meer et al., 2024), exhaustiveness (Khosravani et al., 2024), and coherence (Shan & Lu, 2025)—factors crucial to high-quality argument outputs.
5 Argument Evaluation
Argument evaluation is a higher-level task in argument mining that assesses the quality, strength, and sufficiency of arguments. It encompasses multiple aspects including convincingness, sufficiency, bias, quality ranking, and interpretability. The growing research in this area has led to the development of both theoretical frameworks and empirical methods leveraging machine learning and large language models (LLMs).
5.1 Convincingness Evaluation
Convincingness refers to how persuasive or compelling an argument is. Early studies such as Chalaguine & Schulz (2017) showed that simple linguistic and structural features could suffice for assessing convincingness in online debates. Later, more advanced models were introduced. Gu et al. (2018) explored incorporating topic aspects in evaluating comment convincingness on social platforms, while Potash et al. (2019) proposed ranking mechanisms for argument passages.
The impact of emotions has also been investigated. Chen & Eger (2025) introduced an LLM-based manipulation check system to explore whether emotional language biases human and machine assessments of argument convincingness.
5.2 Sufficiency and Quality
Sufficiency evaluation addresses whether an argument provides enough support for a given claim. A key contribution is Gurcke et al. (2021), who assess sufficiency via conclusion generation, operationalizing the idea that sufficient arguments should enable accurate inference of the claim. Similarly, Liu et al. (2024b) proposed CASA, a causality-aware method for argument sufficiency assessment.
Further formal analysis of sufficiency was provided by Borg & Bex (2024), who explored the concepts of minimality, necessity, and sufficiency in the context of explanation and argumentation. Quality assessment tasks often intersect with sufficiency. For instance, Favreau et al. (2022) applied learning-to-rank BERT models to evaluate argument quality in educational settings.
5.3 Automatic Evaluation and End-to-End Frameworks
Recent works aim to integrate evaluation into end-to-end pipelines. Li et al. (2024b) proposed a multi-task dataset encompassing summarization and evaluation components, enabling automatic scoring of argument relevance and strength. Ruiz-Dolz et al. (2023) introduced a graph-based semantic network to evaluate debate quality using argument graph theory and neural modeling.
A novel coverage metric for summarization tasks was proposed by Khosravani et al. (2024), emphasizing exhaustiveness and accuracy in key point extraction, which reflects evaluation criteria in real-world persuasive writing. From a logic-oriented angle, Burrieza & Yuste-Ginel (2021) examined evaluation through justification logic in multi-agent environments. In education, Winkle et al. (2020) analyzed how teachers apply critical questions to evaluate student arguments, while Rach et al. (2020) benchmarked argument search and evaluation systems in dialogue settings.
5.4 Theoretical and Domain-Specific Perspectives
Beyond purely computational methods, there is growing attention to theoretical and domain-specific evaluation. Palmieri (2024) applied the Argument Mapping Technique (AMT) in corporate controversies to assess argument quality using critical questions. In legal contexts, Grundler et al. (2024) developed AMELIA, an evaluation benchmark for Italian legal documents.
Biases in argument evaluation have also been studied. Diana et al. (2020) investigated adaptive instruction methods to mitigate value-based biases. Baccini & Hartmann (2022) formalized the well-known myside bias using Bayesian models.
6 Future Directions
As argument mining transitions into the era of large language models (LLMs), a wide range of opportunities emerge—alongside challenges that must be systematically addressed. This chapter synthesizes key limitations identified across prior sections and outlines promising future research directions spanning technical, theoretical, and practical dimensions.
6.1 Model Interpretability and Controllability
While LLMs show strong performance across tasks such as stance detection, argument generation, and summarization, their outputs often lack transparency and controllability. In high-stakes domains such as law (Elaraby & Litman, 2022; Xu & Ashley, 2023), healthcare, or political discourse, black-box behavior can hinder trust and adoption.
Future work should prioritize model interpretability, perhaps via structured intermediate representations, symbolic scaffolds (e.g., argument diagrams), or rationale-based generation (Lee et al., 2024). Alignment techniques like self-refinement, chain-of-thought prompting, or rule-augmented decoding may further improve controllability in argumentative outputs.
6.2 Data Quality, Bias, and Hallucination Risks
Several studies, such as Chen & Eger (2025) and Li et al. (2025), have raised concerns regarding the influence of emotion, ideology, or prompt phrasing on argument quality and stance. LLMs trained on massive web corpora may inherit or amplify biases, leading to unbalanced or misleading arguments.
Moreover, hallucination remains a critical challenge in argument generation, particularly when models are required to retrieve or construct supporting evidence. This calls for the integration of fact verification systems, grounding in structured knowledge, or retrieval-augmented techniques (Roush et al., 2024).
6.3 Generalization and Robustness
Despite notable successes in zero- and few-shot settings, many models struggle to generalize across domains (e.g., legal, social media, political debate). Argument summarization models trained on English legal corpora often fail on dialogic or multilingual data (Zhao et al., 2023). Furthermore, domain shifts and adversarial examples can degrade performance in stance classification or evidence retrieval tasks.
Research should focus on domain adaptation, robustness to noisy or adversarial inputs, and multi-lingual evaluation—perhaps through multi-task, curriculum, or meta-learning frameworks.
6.4 Synthetic Supervision and Benchmark Quality
The rise of LLM-generated datasets (e.g., ZeroStance (Zhao et al., 2024), OpenDebateEvidence (Roush et al., 2024)) introduces a double-edged sword. While they enable rapid scaling and coverage, they risk data leakage (pretraining contamination), superficial reasoning, or poor coverage of edge cases.
Future efforts should investigate hybrid data curation pipelines that blend human oversight, synthetic diversity control, and task-specific quality metrics (Khosravani et al., 2024). Benchmarking standards need to evolve accordingly, with support for open-ended outputs, argument graphs, or quality dimensions like sufficiency, diversity, and persuasiveness.
6.5 Integration with Argumentation Theory
Despite decades of theoretical work on argumentation (e.g., Toulmin models, Dung frameworks), most LLM-based methods remain disconnected from formal theory. Argument sufficiency, support/attack structure, or minimality principles are rarely encoded into training objectives or evaluation metrics.
Bridging this gap may involve using theory-guided prompts, symbolic constraints, or structure-aware generation and evaluation systems (Borg & Bex, 2024; Palmieri, 2024). Such alignment could improve model reasoning, transparency, and downstream decision support.
6.6 Human-Centered and Interactive Argument Mining
Finally, the shift toward interactive systems—e.g., debate coaching bots, civic deliberation platforms, or educational feedback tools—demands models that support real-time, multi-turn, and user-aligned argument mining. This includes tracking dialog acts, modeling myside bias (Baccini & Hartmann, 2022), and adapting feedback to user values or learning goals (Diana et al., 2020).
Future systems must integrate HCI design, personalization strategies, and continual learning mechanisms to realize the full societal potential of LLM-based argumentation systems.
7 Conclusion
In this survey, we reviewed the evolution of argument mining in the era of large language models (LLMs), outlining the key tasks—ranging from claim detection and evidence retrieval to stance classification, argument summarization, and evaluation. We categorized modeling approaches before and after the LLM shift, highlighting the increasing role of prompt-based, generative, and retrieval-augmented techniques.
While LLMs have significantly advanced performance across argument mining tasks, they also introduce new challenges related to interpretability, robustness, data quality, and alignment with argumentation theory. As research moves forward, we emphasize the need for better grounding, human-centric interaction, and theoretically informed evaluation.
We hope this survey offers a comprehensive reference for scholars and practitioners seeking to build more reliable, transparent, and cognitively aligned systems for argument understanding and generation.