### v3

# Large Language Models in Argument Mining: A Survey

Abstract
Argument Mining (AM), a critical subfield of Natural Language Processing (NLP), focuses on extracting argumentative structures from text. The advent of Large Language Models (LLMs) has profoundly transformed AM, enabling advanced in-context learning, prompt-based generation, and robust cross-domain adaptability. This survey systematically synthesizes recent advancements in LLM-driven AM. We provide a concise review of foundational theories and annotation frameworks, alongside a meticulously curated catalog of datasets. A key contribution is our comprehensive taxonomy of AM subtasks, elucidating how contemporary LLM techniques—such as prompting, chain-of-thought reasoning, and retrieval augmentation—have reconfigured their execution. We further detail current LLM architectures and methodologies, critically assess evaluation practices, and delineate pivotal challenges including long-context reasoning, interpretability, and annotation bottlenecks. Conclusively, we highlight emerging trends and propose a forward-looking research agenda for LLM-based computational argumentation, aiming to strategically guide researchers in this rapidly evolving domain.
1 Introduction
Argument Mining (AM) is a fast-growing subfield of Natural Language Processing (NLP) concerned with the automatic identification and extraction of argumentative structures—claims, premises, and the relations between them—from textual discourse (Lawrence and Reed, 2019; Patel, 2024). Such structure is central to understanding not just what opinions are expressed but why, enabling applications ranging from debate moderation and fact checking to legal analytics and scientific discourse analysis. Indeed, by transforming unstructured text into structured argument representations, AM supports reasoning-centered applications that go far beyond traditional sentiment analysis, which can determine opinions but not their underlying rationale (Lawrence and Reed, 2019).
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio and Villata, 2018). Foundational surveys—most notably those by Lawrence and Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence and Reed, 2019); Peldszus and Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus and Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
Building on these foundations, subsequent studies have enriched the field with domain- and method-specific insights. Cabrio and Villata (2018) offered a data-driven reflection on the field’s early corpus and model evolution , while targeted surveys explored annotation practices, tool ecosystems, and argument quality assessment. For example, Patel (2024) traces the integration of traditional machine learning with emerging neural techniques; Schaefer and Stede (2021) focus on Twitter; Hua et al. (2019) and Fromm et al. (2021) examine peer-review text; and Guida et al. (2025) assess the performance of LLMs on enforcing topic-specific argument tasks in online comments (Altemeyer et al., 2025; Chen et al., 2024).
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, examining how advances in one area reshape assumptions in another, and reconnects these technical developments to foundational argumentation theory and annotation practice. We treat argument mining as an interlocking system whose components—tasks, data, metrics, and models—evolve together under the pressure of foundation-model capabilities.
This survey aims to fill that gap. We treat argument mining as an interlocking system whose components, tasks, data, metrics, and models, evolve together under the pressure of foundation-model capabilities. Our review makes three contributions: (i) Task re-appraisal. Chapter 3 revisits the canonical AM sub-tasks—claim and evidence detection, relation parsing, stance, argument quality, and summarisation showing how prompting, chain-of-thought, and retrieval-augmented generation blur established boundaries. (ii) Data and evaluation synthesis. Chapters 4 and 5 map the co-evolution of datasets and metrics, highlighting how LLM-assisted annotation and soft-label evaluation change what “state-of-the-art” means. (iii) Technology survey and agenda-setting. Chapter 6 distils common architectural patterns for LLM-era AM (instruction tuning, role-prompt cascades, graph-augmented decoders), while Chapter 7 identifies pressing open problems: long-context reasoning, multimodal and multilingual robustness, bias auditing for machine-generated data, and cost-efficient deployment.
2 Methodology
To obtain a comprehensive, up-to-date picture of research on argument mining (AM) in the large-language-model era, we followed a three-step protocol inspired by prior survey work (Ivanova et al., 2024). Figure 1 gives an overview.
Sources. We first manually queried two bibliographic repositories that jointly cover virtually all peer-review venues in computational linguistics and artificial-intelligence research: (i) the Digital Bibliography and Library Project [^1] https://dblp.org (DBLP) and (ii) the ACL Anthology [^2] https://aclanthology.org. Because the terminology used by the field is heterogeneous, we constructed a Boolean query that targets both generic and fine-grained AM vocabulary (e.g. argument, claim, premise, evidence, key point) and action like mining, extraction, detection. Following recent survey conventions, we limited the search to the last ten calendar years (2015–2025, inclusive), specific for recent 4 years of LLMs era. We find around 210 unique publications that match the query and fall into the 10-year window. We de-duplicated pre-prints against their camera-ready versions by DOI.
Dataset
Many AM papers introduce new corpora or benchmarks that subsequently shape the evaluation landscape. We therefore flagged all items whose title or booktitle fields contained any of the strings dataset, corpus, benchmark, collection, or shared task. This yielded around 70 dataset papers, i. e., works that either (a) release an original annotated resource or (b) substantially extend an existing one. In this work we focus on 40 of represent work to discuss.
Iterative Snowballing
To mitigate false negatives, especially papers that use idiosyncratic task names (e.g. “key-point analysis”, “argument quality ranking”) we employed one rounds of backward and forward snowballing on the seed set’s references. The procedure added 38 further works (6.9 % of the corpus), all of which passed the same year and scope filters. Because every snowballed item also appears in DBLP or the ACL Anthology, the final corpus size is 250 research papers.
Figure 1: Paper-collection workflow and resulting dataset.
3 Task Taxonomy
3.1 Claim Detection
Claim detection identifies central propositions requiring justification across diverse textual landscapes. This task forms the cornerstone of argument mining pipelines, with methodologies evolving to address domain-specific complexities and data modalities. The research landscape spans multiple dimensions:
3.1.1 Platform-Specific Detection Environments
Social Media & Web Platforms: These environments demand specialized approaches for informal, dynamic content. Beltrán et al. (2021) developed ClaimHunter for real-time Twitter monitoring, while Berendt et al. (2021) created FactRank for Dutch-language verification. Multimodal integration significantly enhances detection capabilities, as demonstrated by Cheema et al. (2022) through their MM-Claims dataset combining visual and textual evidence. The CLAIMSCAN framework Sundriyal et al. (2023) further advances social media truth discovery through comprehensive claim span identification.
Video & Multimodal Environments: Emerging research addresses claims in audiovisual content. Rayar (2024) pioneered multimodal video analysis techniques, while Giedemann et al. (2025) contributed ViClaim - a multilingual benchmark dataset specifically designed for video-based claim detection. These approaches must overcome unique challenges including speech-to-text conversion errors and visual-textual semantic alignment.
3.1.2 Domain-Specialized Detection Applications
Healthcare & Biomedical: This critical domain combines fraud detection Gao et al. (2019), privacy-preserving blockchain verification Mohan and Praveen (2019), and evolutionary optimization Tubishat et al. (2025). Biomedical applications focus on Twitter-based health claims Wührl and Klinger (2021) and contradictory research detection Yazi et al. (2021). Recent systematic reviews du Preez et al. (2025) evaluate machine learning efficacy across healthcare fraud patterns.
Environmental Claims: Targeting greenwashing and sustainability misinformation, researchers have developed specialized detectors Woloszyn et al. (2021) and comprehensive frameworks Stammbach et al. (2023). These systems must navigate scientific terminology and corporate reporting nuances while identifying substantiated environmental assertions.
Financial Sector: Financial claim detection addresses numerical assertions and market misinformation through systems like NumClaim Chen et al. (2020) for investor-oriented analysis. Novel datasets Shah et al. (2024) enable weak-supervised detection of numerical claims in earnings reports and market analyses, addressing domain-specific challenges like quantitative reasoning.
3.1.3 Contextual Application Scenarios
Educational Contexts: Claim detection enhances academic writing assessment and scientific discourse analysis. Wan et al. (2020) established correlations between claim presence and writing quality, while Hafid (2024) developed interpretation frameworks for scientific claims in online discourse, enabling more effective knowledge extraction from academic conversations.
Political Fact-Checking: This high-stakes domain employs innovative fairness techniques like actor masking Dayanik and Padó (2020) and positive unlabeled learning Wright and Augenstein (2020). Advanced verification frameworks Jafari and Allan (2024) improve robustness against adversarial political messaging, with rumor detection systems Lin et al. (2021) using claim-guided hierarchical attention networks to combat misinformation.
3.1.4 Cross-Lingual Methodologies
Multilingual detection addresses global misinformation through adapter fusion techniques Schlicht et al. (2023) and LLM-based approaches Vykopal et al. (2025). Cross-lingual transfer learning Panchendrarajan and Zubiaga (2025) enables knowledge sharing between languages, while transformer-based multilingual fact claiming Böck et al. (2021) demonstrates effectiveness across diverse linguistic landscapes.
3.1.5 Verification and Anomaly Detection
Claim Verification: Sophisticated validation frameworks Irnawan et al. (2025) assess claim authenticity through evidence retrieval and reasoning chains. Robust verification techniques Jafari and Allan (2024) address adversarial examples and out-of-distribution claims in high-risk domains.
Anomaly Detection: Specialized techniques identify fraudulent patterns through uncertainty modeling Vandervorst et al. (2024) and profile-based feature engineering Duval et al. (2022). These approaches analyze claim distributions across temporal, behavioral, and semantic dimensions to flag suspicious assertions in insurance, healthcare, and financial domains.
3.2 Evidence Detection
Evidence detection, a critical sub-task within argument mining and misinformation identification, involves identifying textual or multimodal components that substantiate or refute a particular claim. Recent advances in this area include diverse approaches ranging from leveraging multimodal evidence to employing graph-based methods and pretrained language models.
3.2.1 LLMs in Evidence Detection
Several studies have focused on employing pretrained language models for automated evidence extraction. Dementieva and Panchenko Dementieva and Panchenko (2021, 2020) demonstrated the effectiveness of cross-lingual evidence in bolstering monolingual fake news detection. Elaraby and Litman Elaraby and Litman (2021) employed self-trained pretrained language models specifically optimized for evidence detection, achieving notable improvements. Graph-based techniques also constitute a vital approach within evidence detection. Xu et al. Xu et al. (2022a, b) introduced an evidence-aware fake news detection method utilizing graph neural networks to mine fine-grained semantics. Moreover, Guo et al. Guo et al. (2023) developed an interpretable fake news detection model leveraging graph-based evidence, enhancing interpretability and accuracy. Interactive and dynamic evidence retrieval methodologies have also emerged. Liao et al. Liao et al. (2023a, b) proposed a multi-step evidence retrieval enhancement framework (MUSER), significantly improving accuracy by iterative evidence refinement. Furthermore, Wu et al. Wu et al. (2024) developed a unified evidence enhancement inference framework that dynamically adapts to incoming information.
3.2.2 Multilingual Evidence Detection
A significant trend involves evidence extraction across multiple sources or languages. Dementieva et al. Dementieva et al. (2023) proposed a multilingual approach named Multiverse, integrating evidence from diverse linguistic contexts. Hammouchi and Ghogho Hammouchi and Ghogho (2022) similarly developed an evidence-aware multilingual fake news detection system, underscoring the importance of linguistic diversity in robust misinformation detection.
3.2.3 Cross-domain Evidence Detection
Cross-domain adaptability and out-of-distribution scenarios have also garnered attention. Liu et al. Liu et al. (2024a) addressed out-of-distribution issues through dual adversarial debiasing, ensuring evidence-aware models remain effective when encountering novel misinformation types. Additionally, Ferdush et al. Ferdush et al. (2025) focused on cross-domain fake news detection by integrating evidence from multiple social media platforms, enhancing cross-platform robustness.
3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu and Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
3.2.5 Counter-Evidence Detection
Evidence detection is also critical for identifying contradictions and managing conflicts in factual claims. Jiayang et al. Jiayang et al. (2024) introduced the ECON framework for detecting and resolving evidence conflicts, an essential capability for maintaining informational integrity. Kharrat et al. Kharrat et al. (2022) further proposed a contradiction detection approach utilizing semantic relations and uncertainty evidence. Explainability remains integral to the credibility of evidence detection systems. Dong et al. Dong et al. (2024) proposed EMIF, an evidence-aware multi-source information fusion network emphasizing explainability, critical for transparency and trust in automated misinformation detection.
3.3 Relation Identification
If claim detection marks the nodes of an argumentative graph, relation identification supplies its edges. Early work treated the task as sentence-pair classification—“support”, “attack”, or “unrelated”—in tightly controlled genres such as German news commentary (Eckle-Kohler et al., 2015). Genre expansion quickly revealed edge sparsity and annotation drift: the AMPERSAND corpus of Reddit threads, for example, required additional labels like rebuttal and mitigation to capture conversational nuance (Chakrabarty et al., 2020). Methodologically, the field has moved from feature-rich SVMs to graph-neural encoders that propagate information across discourse trees (DBLP:conf/acl/ZhangNL23). Yet even state-of-the-art models stumble when relations span paragraphs, rely on implicit premises, or cross language boundaries. .
3.4 Argument-Quality Assessment
Beyond structural edges lies the question of how good an argument actually is. Early datasets split quality into separate facets—logical sufficiency (Stab and Gurevych, 2017), rhetorical convincingness (Habernal and Gurevych, 2016), or holistic five-point scales (Lauscher et al., 2020). Recent work combines these dimensions and asks models to assign soft labels that reflect annotator disagreement (Plank, 2022). Large language models have further blurred boundaries: prompt-based “LLM judges” now produce zero-shot sufficiency scores that rival supervised baselines, but also inherit topic and sentiment biases (Chen and Eger, 2025). Despite impressive headline gains, fine-grained error analyses show that LLMs still favour fluent but logically thin arguments over less polished yet better-supported ones.
3.5 Stance Detection
Stance detection is the task of automatically identifying an author’s position towards a specific target, claim, or topic — typically labeled as pro, con, or neutral. This task plays a critical role in argument mining, especially in opinionated domains such as social media, online debates, and political discourse.
Datasets and Evaluation.
Recent efforts have produced a variety of stance detection datasets to support research in this area. For instance, Niu et al. (2025a, b) introduce C-MTCSD, a Chinese multi-turn conversational stance detection dataset, which captures nuanced stance evolution. Gyawali et al. (2024) contribute GunStance, a dataset focusing on gun regulation debates. These datasets facilitate stance classification in diverse linguistic, political, and conversational contexts.
Classical and Neural Methods.
Early neural approaches, such as those in Hosseinia et al. (2019), leverage BiLSTM architectures to detect stance in argumentative opinions. More recent work like Wang et al. (2024) introduces a recurrent stick-breaking topic model to jointly model stance and topic clusters. Furthermore, approaches like Li et al. (2024c) simulate rumor propagation to enhance stance prediction under social influence.
Large Language Models for Stance Detection.
With the rise of LLMs, several works have evaluated or leveraged their capabilities. Zhao et al. (2024) propose ZeroStance, using ChatGPT to generate datasets and perform zero-shot stance detection. Gambini et al. (2024) conduct a systematic evaluation of various LLMs (e.g., GPT-3, BLOOM) for stance detection on social media. Complementary to this, Lee et al. (2024) introduce LOGIC, where LLM-generated rationales guide small models to improve stance reasoning.
Cross-lingual and Domain Adaptation.
Multilingual stance detection remains a challenge. Charfi et al. (2024) present a multi-dialect Arabic stance corpus and cross-domain benchmarks. Works like Alghaslan and Almutairy (2024); Shukla et al. (2024) leverage LLM fine-tuning and ensembling for Arabic stance detection tasks.
Zero-shot and Few-shot Generalization.
The promise of zero-shot stance detection has spurred further innovation. Ma et al. (2024a, b) propose Chain-of-Stance, where LLMs perform chain-of-thought style reasoning. Guo et al. (2024) incorporate structured external knowledge from LLMs to improve generalization. Zhang et al. (2024) integrate commonsense knowledge through adversarial learning.
Bias Mitigation and Calibration.
Acknowledging LLM bias, Li et al. (2025, 2024a) propose counterfactual augmented calibration strategies to reduce stance-related biases in LLM outputs.
Social Media and Health Applications.
Beyond generic domains, stance detection has been used in medical and health misinformation contexts. Davydova et al. (2024) analyze COVID-19 tweet stances; Sun et al. (2024a) study HPV vaccine stance detection using in-context learning and fine-tuning. Akash et al. (2024) assess the feasibility of open-target stance detection using LLMs.
Synthetic and LLM-generated Data.
LLM-generated synthetic datasets are also explored. For example, Wagner et al. (2025) show that LLM-synthesized stance data can outperform traditional fine-tuning when training low-resource classifiers. Similarly, Fraile-Hernandez and Peñas (2024) demonstrate the use of generative LLMs in political activism scenarios.
Language-independent Embeddings.
Finally, de Landa and Agerri (2025) present language-agnostic methods using interaction-based embeddings and multilingual LLMs to promote cross-lingual stance transfer.
In sum, the task of stance detection has evolved from simple text classification to a rich and multi-faceted research area, with contributions spanning LLM prompting, data generation, bias handling, and multilingual adaptability.
3.6 Argument Summarization
Argument summarization focuses on generating concise and informative summaries of argumentative texts, including debates, discussions, legal decisions, and online forums. Compared to general summarization, this task presents unique challenges due to the complexity of argumentative structure, the presence of contrasting stances, and the need to preserve key claims, evidence, and discourse relations.
Early work such as ConvoSumm(Fabbri et al., 2021) laid the foundation for integrating argument mining with abstractive summarization in dialogues, demonstrating the benefits of identifying argumentative components in conversation summarization. Building on this, several works have extended the scope of argument summarization to new domains such as law(Elaraby et al., 2023; Elaraby and Litman, 2022; Xu and Ashley, 2023; Steffes and Rataj, 2022), where legal documents are often long and require nuanced treatment of argumentative structure. The ArgLegalSumm model(Elaraby and Litman, 2022) proposed incorporating argument mining to enhance legal document summaries, and was further extended with reranking modules to improve informativeness and coherence(Elaraby et al., 2023).
In terms of modeling techniques, Shan and Lu(Shan and Lu, 2025) applied multi-task learning with argument relation-aware contrastive learning to better capture opinion divergence in debates. Similarly, Syed et al.(Syed et al., 2023) introduced frame-oriented summarization to capture pragmatic aspects of arguments in discussions. Other efforts, such as by Meer et al.(van der Meer et al., 2024), analyzed the diversity of generated summaries, revealing trade-offs between informativeness and representation of varied perspectives.
Dataset construction remains critical for progress. OpenDebateEvidence(Roush et al., 2024) provides a massive-scale benchmark with paired argument and evidence annotations, enabling more robust training. Likewise, ORCHID(Zhao et al., 2023) and ConvoSumm(Fabbri et al., 2021) support summarization across argumentative dialogues in Chinese and English, respectively.
3.6.1 Key Point Analysis
Key Point Analysis (KPA) is a fine-grained subtask within argument summarization that focuses on identifying representative and recurring ideas—referred to as “key points”—from a collection of arguments. Introduced by Bar-Haim et al. (Bar-Haim et al., 2020), KPA provides a structured framework for summarizing public opinions and argumentative discourse in a more interpretable and scalable manner than free-form summarization.
Initial approaches relied on clustering or semantic similarity. For example, Bar-Haim et al.(Bar-Haim et al., 2020, 2021) introduced the idea of extracting key points and matching arguments to them using supervised and unsupervised approaches, which inspired multiple shared tasks. Phan et al.(Phan et al., 2021) proposed a simple yet highly accurate model using bi-encoders for key point matching, while Reimer et al. (Reimer et al., 2021) explored pretrained encoders and alignment strategies.
Later efforts integrated contrastive learning (Alshomary et al., 2021) and graph-based representations (Li et al., 2024d) to improve robustness and scalability. Incremental and real-world-ready systems were also developed, as seen in Eden et al. (Eden et al., 2023).
With the rise of LLMs, researchers explored their generative and representational capabilities. Tang et al.(Tang et al., 2024b, a) examined aspect-based and prompted KPA for review summarization, highlighting the flexibility of LLMs in domain adaptation. Li et al.(Li et al., 2023) proposed an iterative clustering and abstractive summarization framework to automatically induce key points. More recently, Fukuma et al. (Fukuma et al., 2024) demonstrated that zero-shot KPA is feasible through LLM-based question generation and network analysis.
Overall, KPA bridges summarization, clustering, and representation learning. It has proven effective in condensing opinion-rich texts and continues to be enriched by LLM capabilities and multi-step reasoning frameworks.
4 Dataset and Benchmark
Dataset Year Size Lang. Modality Main Tasks LLM Origin33footnotemark: 3 Open Source
ArgMinNews 2015 80 DE Text Argument Relation Identification
UKP OpposingArgumentsInEssays 2016 402 EN Text Claim Detection
UKP ConvArg 2016 16k EN Text Convincingness Checking
UKP InsufficientArguments 2017 1 k EN Text Sufficiency Checking
IBM Claim Stance Dataset 2017 2.3k EN Text Stance Detection
IBM Claim Sentences Search 2018 1.49M EN Text Claim Detection
IBM Evidence Quality 2019 5.6K EN Text Claim Detection
PERSPECTRUM 2019 20k EN Text Claim Detection & Evidence Detection
UKP ASPECT 2019 3.5K EN Text Claim Detection
IBM Labeled Emphasized Words in Speech 2018 2.4k EN Multi Text to Speech
AMPERSAND 2020 2.7k EN Text Argument Relation Identification
CoPA 2019 800 EN Text Claim Detection
IBM ArgQ-14kArgs 2019 14k EN Text Convincingness Checking
IBM Argumentative Sentences in Recorded Debates 2020 700 EN Text Claim Detection
Review-Rebuttal 2020 4.7K EN Text Claim Detection
IBM Evidences Sentences 2020 29k EN Text Convincingness Checking
IBM XArgMining 2020 35k DE,NL,E S,FR,IT Text Claim Detection & Evidence Detection & Stance Detection
IBM Recorded Debating Dataset 2020 3.5k EN Multi Speech to Text
MEDLINE 2020 6.7k EN Text Claim Detection & Argument Relation Identification
GAQCorpus 2020 5k EN Text Convincingness Checking & Sufficiency Checking
IBM ArgKP 2021 27k EN Text Key Point Analysis
SciARK 2021 1k EN Text Claim Detection & Evidence Detection
IAM 2022 69k EN Text Claim Detection & Evidence Detection & Stance Detection
FinArg 2022 136 EN Text Claim Detection & Stance Detection
Sustainable Diet Arguments on Twitter 2022 592 EN Text Claim Detection & Evidence Detection & Stance Detection
Modality: text, audio, video, multi. ‘LLM Origin’ ✓ = created with or for LLM evaluation.
Table 1: Recent corpora; shaded rows belong to the “LLM era” (2021–present). German (DE), Dutch (NL), Spanish (ES), French (FR), and Italian (IT).
Dataset Year Size Lang. Modality Main Tasks LLM Origin44footnotemark: 4 Open Source
Kialo 2022 310k EN Text Argument Relation Identification
IBM SurveyKP 2023 15k EN Text Key Point Analysis
ArguGPT 2023a 8k EN Text Claim Detection
StoryARG 2023 2.4k EN Text Claim Detection & Argument Relation Identification
GENEVA 2023 3k EN Text Argument Relation Identification
VivesDebate-Speech 2023 29 EN Multi Speech to Text
ArgSciChat 2023 498 EN Text Argument Relation Identification
MuLMS-AZ 2023 10k EN Text Argument Relation Identification
ArgAnalysis35K 2023 35k EN Text Argument Relation Identification
ORCHID 2023 14k EN Text Argument Summarization & Stance Detection
ARIES 2024 81k EN Text Argument Relation Identification
FS150T Corpus 2024 21k EN Text Claim Detection & Stance Detection
CasiMedicos-Arg 2024 7k EN Text Claim Detection & Stance Detection
CEAMC 2024 226 EN,ZH Text Claim Detection & Evidence Detection & Stance Detection
ASE 2024b 2.3 k EN Text Evidence Detection & Convincingness Checking & Argument Summarization & Sufficiency Checking
ConQRet 2025 6.5k EN Text Evidence Detection & Argument Quality Assessment
Modality: text, audio, video, multi. ‘LLM Origin’ ✓ = created with or for LLM evaluation.
Table 2: Continue from previous table. Recent corpora belong to the “LLM era” (2021–present). German (DE), Dutch (NL), Spanish (ES), French (FR), and Italian (IT). Chinese (ZH)
Annotated corpora have always served a dual function in argument–mining research: they are the evidence on which empirical claims are validated, and they are the scaffolding that decides which problems can be modelled at all. Historically, dataset design followed a “task-first” logic—each new sub-task (claim detection, convincingness ranking, sufficiency checking, key-point analysis) first appeared as a corpus release and only thereafter as a stream of modelling papers. Compared with the small, task-specific, English-only corpora that served feature-based models, today’s “LLM-era” datasets are larger, multi-label, cross-domain, and often partly machine-generated—properties chosen explicitly to match the breadth, multitask capacity, and data appetite of modern instruction-tuned language models. In this work, we conduct a comprehensive survey of how datasets and benchmarks meet the requirements and adapt to evolution of the LLM era.
Specifically, we outline it in three successive research phases: (i) from task-specific, human-curated collections to integrated, LLM-centric resource (Section 4.1), (ii) Using LLM as annotator and generator (Section 4.2, and (iii) an emerging focus on scale, modality, and language breadth (Section 4.3). We review each phase and close look at over 40 datasets (shown in Table 1 and 2 to figure out the trends and open challenges.
4.1 From One Task–One Dataset to Integrated Corpora
[…]
Single-purpose design proved invaluable for methodological clarity. Because each dataset held only one variable constant, researchers could dissect feature contributions and model architectures with minimal confounds. Notably, margin-ranking SVMs on ConvArg delivered a 19-point Pairwise Accuracy gain over bag-of-words in large part because the dataset’s simplicity isolated ranking logic from component detection errors (Habernal and Gurevych, 2016). However, this fragmentation soon impeded cumulative progress. Cross-task transfer became almost impossible: a model trained to find claims in UKP essays failed on biomedical abstracts, not because the notion of a claim differed but because no shared label inventory or evaluation split existed across corpora (Mayer et al., 2020). Moreover, meta-analysis was thwarted by incompatible annotation units (sentence vs. clause) and divergent definitions of “support” and “attack” (Reimers et al., 2019).
Faced with these limitations—and armed with instruction-tuned large language models capable of multi-task learning—dataset designers began to integrate multiple label layers within a single corpus. The turning point was IAM (69 k Reddit threads), which simultaneously marks clause boundaries, supporting evidence links, stance polarity, and topic metadata. Multi-task RoBERTa trained on IAM improved macro-F1 by 4–6 points over single-task baselines on both claim detection and stance classification, demonstrating that shared representations benefit from joint supervision (Cheng et al., 2022). Similarly, ARIES (81 k documents) overlays RST-style discourse graphs on argumentative zones; graph neural networks trained here transfer to the cross-lingual MuLMS-AZ news corpus with only a 5-point F1 loss, far less than the 15-point drops reported for earlier single-task systems (Gemechu et al., 2024; Schrader et al., 2023). Integrated corpora are not confined to component-level annotation: ORCHID fuses stance, key-point matching, and abstractive summaries in 14 k opinion paragraphs, enabling the first evaluation of summarisation metrics grounded in argumentative structure (Zhao et al., 2023); StoryARG does likewise for narrative argumentation (Falk and Lapesa, 2023).
[…]
Research gaps and future work.
(1) Reliability audits: integrated corpora should report per-label agreement and analyse cross-label interference (e.g., does stance bias sufficiency?). (2) Shared ontologies: without a common inventory, the promise of multi-task modelling will stall at dataset borders. (3) Task balancing: early evidence suggests that over-represented labels dominate learning; dynamic sampling or curriculum learning could mitigate this skew. Addressing these issues will turn integrated corpora from mere “data lakes” into coherent, theory-driven benchmarks suitable for the next generation of argument-aware language models.
4.2 LLM-Assisted Creation and Annotation
Large language models are rapidly becoming collaborators in corpus construction rather than merely downstream consumers. Their involvement takes three main forms—prompt-driven synthesis, machine-in-the-loop annotation, and fully automatic labelling of difficult variables—and each reshapes long-standing assumptions about cost, bias and reliability.
Prompt-driven synthesis
gained prominence when Liu et al. (2023a) used GPT-4 to draft thousands of claim–counter-claim pairs across diverse topics. Although human editors rejected a non-trivial minority for logical incoherence, the retained material exhibits argumentative structures broad enough to boost zero-shot performance on earlier convincingness benchmarks by a noticeable margin—often up to ten percent relative improvement. Stylistic audits, however, reveal a homogenised, over-polite register and a marked reduction in hedging compared with organically written debate, raising concerns that synthetic data may narrow the stylistic bandwidth of subsequent models (Long et al., 2024).
Machine-in-the-loop annotation
Humans remain in control but delegate first-pass labelling to an LLM. The ASE corpus exemplifies this workflow: GPT-4 proposes evidence spans, sufficiency judgements and short summaries, after which human reviewers confirm or amend the suggestions (Li et al., 2024b). Verification time falls markedly, yet error analysis shows that chain-of-thought prompting, while improving recall, doubles the rate of subtle hallucinations. Similar hybrid gains are reported for SurveyKP, where key-point clusters refined by annotators achieve noticeably higher purity than crowd-only baselines at roughly half the labour hours (Eden et al., 2023). The flip side is automation anchoring: reviewers tend to accept stylistic quirks or soft politeness biases introduced by the model, an effect already quantified in appropriateness studies by Ziegenbein et al. (2023).
Automatic Judge
A third trend outsources to automatic judgements LLMs when labels are cognitively heavy or prone to low consensus. In ConQRet, GPT-4 decides whether retrieved passages “fully answer” complex questions before experts handle only borderline cases (Dhole et al., 2025). Agreement with specialists is high overall, yet topic-sensitive analyses show the model penalises rhetorical questions more harshly than humans—a bias echoing manipulation-check findings on emotional language by Chen and Eger (2025). Moreover, when the same model family later evaluates systems trained on these labels, circularity becomes a real possibility: sufficiency scores judged by GPT-4 appear several points higher when the evaluator is also GPT-4 than when a conventional RoBERTa baseline is used (Liu et al., 2024b).
Research Gaps and Future Works
These developments create new methodological obligations. Few corpora publish detailed “data cards” that document prompt templates, sampling temperatures, rejection rates or post-editing guidelines, leaving provenance opaque and bias assessment difficult. Systematic audits are needed to understand how demographic, topical and stylistic biases propagate from model to dataset. Equally pressing is the design of adversarial test suites: models trained solely on clean, LLM-generated text degrade sharply on noisy social-media arguments, suggesting that synthetic and natural data must be blended with care. Finally, the legal status of machine-generated content remains unsettled; future releases should clarify licensing to safeguard downstream reuse.
In sum, LLM assistance lowers annotation cost and widens the range of available labels, yet it also imports model-specific artefacts and raises questions about bias, provenance and evaluation circularity. Addressing these issues through transparent documentation, rigorous bias audits and adversarial evaluation will be vital if LLM-centred datasets are to serve as reliable foundations for the next generation of argument-mining models.
4.3 Scale, Modality, and Language
As argument–mining goals have broadened, corpus builders have sought breadth in three orthogonal directions—size, modality, and language. Each dimension has expanded unevenly, yielding fresh modelling opportunities but also exposing new blind spots.
Size: from thousands to hundreds of thousands.
The median pre-LLM corpus contained only a few thousand annotated instances, limiting the depth of neural fine-tuning. Integrated resources now appear at an order-of-magnitude larger scale. Debate platform dumps such as Kialo  [^5] https://www.kialo.com/, with well over a quarter-million crowd-moderated nodes, allow graph-level experiments in argumentation theory that were inconceivable when researchers relied on essay corpora (Agarwal et al., 2022). Large scale is not merely a matter of quantity: it supports adversarial splitting, more granular error analysis, and realistic low-resource simulations. Yet volume can disguise annotation drift; agreement figures reported for the first 10
Modality: beyond plain text.
Argumentative discourse is rarely confined to prose; tone of voice, slide layouts, or imagery all shape persuasiveness. Two speech–text corpora IBM’s Recorded Debating and the VivesDebate-Speech set align transcribed turns with audio and meta-data, enabling first forays into prosody-aware argument strength (Orbach et al., 2020; Ruiz-Dolz and Sanchez, 2023). The Image-Argument pilot goes further, annotating visual claims and layout cues in news pages (Liu et al., 2023b). Early multimodal baselines show detectable gains when acoustic or layout features are fused with text, but they also reveal new annotation pain-points: irony carried by image–caption incongruence or persuasive pauses in speech remain largely uncaptured. More importantly, evaluation metrics such as “pairwise convincingness” have yet to be re-interpreted for cross-modal signals.
Language: cautious steps toward multilinguality.
English continues to dominate, but a second tier of languages,German, Dutch, Spanish, French, Italian, and Chinese, has emerged. Parallel corpora such as XArgMining demonstrate that simple cross-lingual projection recovers much, but not all, of English performance (Toledo-Ronen et al., 2020). Specialised bilingual resources add depth: FinArg pairs English and domain-specific finance jargon with argumentative structure (Alhamzeh et al., 2022), whereas CEAMC offers Chinese–English medical arguments linked by explicit claim–evidence pairs (Ren et al., 2024). Nevertheless, most multi-task corpora remain monolingual, so joint modelling of component, stance and quality across scripts is still aspirational. Preliminary cross-lingual experiments suggest that stance and sufficiency transfer degrade much faster than claim segmentation, hinting that cultural framing affects higher-order argumentative facets more than low-level span boundaries.
Research gaps and future work.
The three expansion axes seldom intersect in the same dataset: no existing corpus is simultaneously large-scale, multimodal, and multilingual. This fragmentation hampers holistic robustness tests—for instance, a model trained on English speech debates cannot yet be evaluated on Chinese medical videos. Future benchmarks should therefore prioritise cross-dimensional design, even if that means curating smaller but more diverse seed sets. Equally urgent is the development of evaluation metrics that remain meaningful when text, audio, and visuals interact. Without such metrics, gains reported on one modality may mask losses in another, leading to brittle real-world systems.
In short, corpora have grown larger, richer, and more linguistically varied, but their growth resembles a patchwork: impressive in individual squares, incomplete as a whole. Closing these gaps will be essential if argument mining is to generalise beyond the Anglophone, text-only laboratory into the multimodal, multilingual public sphere where real arguments live and evolve.
5 Argument Quality Assessment
Figure 2: Overview of quality dimensions for argument mining quality assessment discovered in the surveyed literature. Note: The figure is an extension of a taxonomy proposed in Ivanova et al. (2024)
Argument Quality Assessment (AQA)—also termed Argument Evaluation or Argumentation Quality (AQ), is the stage of the argument-mining pipeline that quantifies how well an already-identified argument fulfils its inferential and communicative objectives. Whereas argument component identification and relation detection answer the structural questions “what is an argument and how is it built?”, AQA addresses the normative question “how good is this argument?”. Early philosophical (Wenzel, 1990) syntheses distinguish three overarching dimensions, Logic/Cogency, Dialectic/Reasonableness, and Rhetoric/Effectiveness, each further decomposed into sub-dimensions such as acceptability, sufficiency, relevance, audience adaptation, or stylistic appropriateness (shown in Figure 2) .
AQA interacts tightly with but is distinct from argument component detection, relation classification, and generation. In retrieval settings, quality scores serve as features to re-rank argumentative passages. In writing-support and educational tools, they trigger targeted feedback on clarity or persuasiveness. Within persuasive dialogue systems, real-time quality estimates guide strategy selection and counter-argument generation. Finally, when large language models (LLMs) act both as generators and judges, AQA becomes a meta-evaluation task, raising new challenges of bias amplification and self-consistency.
Task Definition.
Formally, given a textual unit $a$, ranging from a single premise to a full argumentative thread, and, optionally, contextual information $c$ (topic, dialogue history, background knowledge), AQA seeks to output either (i) a scalar or categorical quality score $q(a,c)\in\mathbb{R}$, (ii) a vector of dimension-specific scores $\mathbf{q}(a,c)$, or (iii) a preference relation $a\succ b$ between pairs $(a,b)$. Approaches thus fall along two orthogonal axes: absolute vs. relative scoring—Likert-style ratings versus pairwise preferences, and intrinsic vs. extrinsic evaluation—judging the argument in isolation versus with external facts or opponent turns . Each design choice trades annotation complexity for construct validity: relative comparisons often raise inter-annotator agreement, yet absolute scales enable downstream aggregation and benchmarking.
The remainder of this section dissects AQA along five research threads: theoretical quality dimensions (Section 5.1), modelling advances from feature engineering to LLM-based judges (Section 5.2), evaluation metrics and meta-evaluation (Section 5.3), and applications (Section 5.4).
Tier Sub-dimension Definition
Logic Acceptability Premises (and claim) do not contradict widely accepted facts or domain conventions (Wachsmuth et al., 2017a).
Local relevance Each premise bears a direct inferential link to its claim; off-topic or redundant information lowers the score (Wachsmuth et al., 2017b).
Sufficiency The collective evidence warrants the claim; absence of missing warrants yields a “fully supported” label (Stab and Gurevych, 2017; Gurcke et al., 2021).
Dialectic Global relevance The argument moves the overall discussion toward issue resolution—e.g., high in IBM Rank-30k “would-you-use” labels (Toledo et al., 2019; Gretz et al., 2020).
Opponent sensitivity Anticipates or rebuts likely objections, showing dialogical awareness (Khatib et al., 2017; Habernal and Gurevych, 2017).
Overall reasonableness Conforms to norms of critical discussion; free of straw man, ad hominem, and similar fallacies (Baff et al., 2018).
Rhetoric Credibility Projected expertise, trustworthiness, and goodwill via source citation, hedging, or professional register (Khatib et al., 2016).
Appropriateness Conformity of tone, politeness, and word choice to social norms and platform guidelines; penalizes toxicity or profanity (Ziegenbein et al., 2023).
Clarity Absence of ambiguity and needless complexity; logical flow and unambiguous wording facilitate comprehension (Persing et al., 2010).
Emotional Appeal Strategic use of sentiment, moral framing, or arousal to motivate the audience (Chen and Eger, 2025).
Arrangement Coherent discourse structure and persuasive ordering of premises, often assessed via discourse markers or paragraphing cues (Persing et al., 2010).
Strength Perceived inherent force or weight of the reasoning, independent of actual attitude change (Lukin et al., 2017).
Convincingness Likelihood that a neutral judge would find the argument compelling in pairwise or listwise comparisons (Gienapp et al., 2020).
Persuasiveness Observed ability to shift audience stance or votes after exposure; measured on platforms such as Change My View (Tan et al., 2016).
Winning Side Debate-level outcome: side whose cumulative arguments secure more post-debate votes or attitude change (Ivanova et al., 2024).
Table 3: 15 core sub-dimensions of argument quality with concise definitions. Parent tiers follow the Logic–Dialectic–Rhetoric triad.
5.1 Quality Dimensions and Theoretical Frameworks
Argument quality is intrinsically multidimensional. During the last decade, two complementary lines of work have shaped the field: (i) theoretical systematisation, rooted in philosophy and argumentation studies (Wachsmuth and Werner, 2020); and (ii) empirical operationalisation, driven by the need to annotate large corpora and to train machine-learning models. We synthesise both strands in three steps.
5.1.1 Classical Tripartite Perspectives: Logic, Dialectic, Rhetoric
Wenzel’s “three perspectives” framework remains the most influential point of departure for computational AQ . It distinguishes a logical (cogency), a dialectical (reasonableness), and a rhetorical (effectiveness) vantage point on every argument (Wenzel, 1990). Wachsmuth et al. (2017a) operationalize the triad for NLP by posing three guiding questions:
  • Cogency Are the premises acceptable, relevant, and sufficient for the claim?
  • Reasonableness Does the argument withstand counterarguments and respond to the target audience’s concerns?
  • Effectiveness Does the discourse persuade its intended audience through credible, emotional, and well-structured presentation?
These perspectives provide the conceptual backbone for nearly all annotation schemes surveyed by Ivanova et al. (2024). The prevalence of the triptych is not merely historical; empirical work shows that annotators perceive and rate these facets with reasonably high reliability when clear instructions are provided (Gretz et al., 2020; Alhamzeh, 2023).
5.1.2 Fine-grained Sub-dimensions
As summarised in Table 3, each macro-perspective of the Logic–Dialectic–Rhetoric triad can be decomposed into finer, operational sub-dimensions. For the logical tier, empirical work most often labels acceptability, local relevance, and sufficiency. The dialectical tier adds discourse-level notions—global relevance, opponent sensitivity, and overall reasonableness. The rhetorical tier widens the lens still further, capturing audience-oriented aspects such as credibility, appropriateness, clarity, emotional appeal, arrangement, and high-level outcome variables (strength, convincingness, persuasiveness, winning side). Early studies showed that even non-experts agree substantially on some facets—most notably sufficiency (Stab and Gurevych, 2017). Yet boundaries remain fluid: “convincingness” blends logical sufficiency with emotional impact, while “persuasiveness” often collapses the full triad into a holistic score. Consequently, recent annotation protocols increasingly employ multi-label or soft-label designs to let a single argument express several quality aspects at once (Plank, 2022).
5.1.3 Comparative Taxonomy Analysis
Taxonomy Year #Fine-grained Dims. Annotated Corpus (size) Notable Features / Extensions
Blair 2011 3 Philosophical foundation of the Logic Dialectic Rhetoric triad.
Wachsmuth et al. 2017a 15 Dagstuhl 15512 ArgQuality ($\approx$300 arguments) First computational taxonomy.
Lauscher et al. 2020 11 GAQCorpus ($\approx$5k arguments) Demonstrates large-scale crowd annotation.
Ivanova et al. 2024 15 Aligns overlapping labels across 211 papers; proposes harmonised quality meta-labels for transfer learning.
Nussbaum et al. 2020 7 98 science-teacher lesson plans Maps Walton’s Critical Questions to classroom feedback; emphasises opponent rebuttal and reasonableness.
Grundler et al. 2024 5 225 Italian legal decisions Adds statute relevance and argument-scheme fit; stresses domain-specific sufficiency and formality.
Table 4: Side-by-side comparison of six argument-quality taxonomies. “Fine-grained dims.” counts leaf-level labels, excluding overall scores.
While the classical triad remains conceptually robust, empirical practice keeps stretching its margins. For computational modelling, the key open question is not whether to refine the taxonomy further, but how to map diverse, possibly subjective sub-dimensions onto unified, explainable representations that large language models can learn and that human assessors can interpret consistently.
As a example, Table 4 shows the changes of such theory framework, contrasts four influential taxonomies: (Blair, 2011), (Wachsmuth et al., 2017a), (Lauscher et al., 2020), (Ivanova et al., 2024) and two domain-specific schemes (legal AQ (Grundler et al., 2024), educational feedback (Nussbaum et al., 2020)). Three trends emerge: (i) Convergence on the tripartite backbone—every scheme retains some form of Logic–Dialectic–Rhetoric. (ii) Progressive granularity—the average number of annotated sub-dimensions grows from 3 (2010) to 11 (2024). (iii) Domain hybridisation new schemes borrow constructs from adjacent fields (fact-checking, toxicity detection, essay scoring), leading to overlapping labels that challenge cross-dataset transfer.
5.2 Modelling Advances: From Feature Engineering to LLM-based Judges
Rule-based Feature Engineering .
Early automatic approaches to argument-quality assessment relied on feature-engineering pipelines. Linear, SVM and gradient-boosting models consumed handcrafted indicators such as length, sentiment, hedge density and discourse connectives, delivering the first reproducible baselines on essays, reviews and Web debates (Habernal and Gurevych, 2017; Wachsmuth et al., 2017a). On Reddit Change My View, simple length and interaction features already yielded reliable convincingness rankings (Chalaguine and Schulz, 2017). Potash et al. were first to cast convincingness as a pairwise ranking problem, introducing a margin-based loss that remains standard today (Potash et al., 2019). Other work encoded topical similarity to improve local relevance (Gu et al., 2018), applied PageRank over argument graphs to estimate global salience (Wachsmuth et al., 2017b) or factual-consistency rules to flag insufficient support (Stab and Gurevych, 2017). Although feature lists were task-specific, they established the feasibility of automatic AQA on modest corpora such as AAE and Webis ArgQuality (Wachsmuth et al., 2017a).
Neural encoders with task-specific heads.
The 2016–2021 period marks an unequivocal shift from sparse inputs to representation learning. Bidirectional LSTMs outperformed bag-of-words on convincingness by capturing sequential cues (Habernal and Gurevych, 2016); CNN–BiLSTM hybrids added early cross-topic generalization on comment threads (Huber et al., 2019). Several groups then moved from classification to ranking losses: margin-based pairwise objectives (Potash et al., 2019) and listwise NDCG targets (Gu et al., 2018) yielded more stable quality orderings. Transformer fine-tuning largely removed feature design: BERT and RoBERTa lifts of 5–10 F1 or $\rho$ on IBM-Rank and GAQ tasks are now well documented (Gretz et al., 2020; Lauscher et al., 2020). Downstream improvements came from metric-learning variants—e.g. Siamese-SBERT with contrastive loss for out-of-domain convincingness (DBLP:conf/argmining/ZhangR23)—and from multi-task heads that jointly predict stance or component type, boosting macro-F1 on low-resource sets (Favreau et al., 2022). Efficient pairwise crowdsourcing accelerated data creation and ranking-loss training (Gienapp et al., 2020), while contextual embeddings enabled intrinsic, reference-free scoring of short arguments (Wachsmuth and Werner, 2020). Together, these advances consolidated neural encoders as the default for AQA.
Knowledge-aware and graph-based architectures.
Sentence-level encoders struggle with logical sufficiency and discourse flow, prompting richer context integration. Gurcke et al. generate candidate conclusions and use textual entailment to operationalise sufficiency (Gurcke et al., 2021). External-knowledge infusion—via retrieval-augmented transformers—raises sufficiency F1 by 3 points on legal AMELIA when factual statutes are injected at inference time (Liu et al., 2024b). Graph neural networks layered on RST or AMT structures improve global relevance, rebuttal recognition, and “winning-side” prediction, notably in legal and scientific debates (Zhang et al., 2023; Galassi et al., 2023). Semantic-network-based rankers outperform sequential baselines on cross-topic challenge sets by explicitly modelling premise-claim-evidence ties (Ruiz-Dolz et al., 2023). Finally, discourse-intrinsic objectives such as the COVERAGE metric capture informativeness without labels and serve as pre-training tasks that transfer to sufficiency and clarity (Khosravani et al., 2024). These structure- and knowledge-aware approaches close much of the gap between surface-form language models and the normative criteria demanded by argumentation theory.
Self-supervised and intrinsic quality objectives.
[…]
LLM-based judges and prompting paradigms.
Large language models have recently moved from feature generators to stand-alone adjudicators. With as few as three in-context exemplars, GPT-4 matches or exceeds fine-tuned RoBERTa on GAQCorpus for pairwise convincingness and sufficiency (Chen et al., 2024). Performance improves further when prompts elicit chain-of-thought rationales and apply self-consistency voting: Wang et al. report a 7-point macro-F1 gain on IBM-Rank after majority-voting across five diverse reasoning traces (Altemeyer et al., 2025). Retrieval-augmented prompting injects domain facts at inference time; on the AMELIA legal benchmark, RAG-GPT-4 lifts sufficiency F1 by an additional 3 points compared with vanilla GPT-4 (Liu et al., 2024b). Beyond accuracy, LLM judges reveal latent biases. Manipulation-check experiments demonstrate that GPT-4 systematically discounts emotionally charged but fallacious premises, whereas earlier transformer baselines over-score them (Chen and Eger, 2025). However, agreement may be inflated when both training data and evaluation sets share topic distributions (Schiller et al., 2022); hybrid pipelines therefore combine frozen LLM judges with lightweight re-rankers or calibration layers to reduce domain drift and latency (Ruosch et al., 2024; Kotelnikov et al., 2022).
5.3 Evolution of Evaluation Metrics
The choice of evaluation metrics has co-evolved with modeling techniques and data regimes in argument–quality assessment (AQA). Over time, three partly overlapping strands have emerged: annotation reliability, task-level prediction quality, and system-level utility. We review each strand chronologically, highlighting how new metrics have responded to changing task formulations and application demands.
From inter-annotator agreement to confidence-weighted labels.
Early corpora with nominal “good / bad” labels reported Cohen’s (Hsu and Field, 2003) or Fleiss’ $\kappa$ (Fleiss et al., 1981). As datasets grew and adopted ordinal or continuous scales, Krippendorff’s $\alpha$ (Krippendorff, 2011) became preferred because it handles mixed measurement levels (Stab and Gurevych, 2017). Large crowdsourced resources such as IBM Rank-30k now accompany raw labels with MACE-style worker weighting; after filtering, $\alpha$ values can reach 0.93, approaching expert reliability (Gretz et al., 2020). A recent meta-survey corroborates these gains across 32 datasets (Ivanova et al., 2024).
Task-level metrics: from categorical F1 to correlation statistics.
When AQA was framed as binary classification, macro-averaged Precision, Recall, and F1 sufficed (Wachsmuth et al., 2017a). As labels became graded, regression metrics—mean absolute error and root MSE—and rank correlations (Pearson (Benesty et al., 2009), Spearman $\rho$ (Hauke and Kossowski, 2011)) took centre stage. Persing & Ng first cast persuasiveness as a six-point regression task (Persing and Ng, 2020), and Lauscher et al. popularised Spearman $\rho$ on GAQCorpus (Lauscher et al., 2020). Multi-task heads that jointly predict stance and several quality facets are usually evaluated with a macro-averaged MAE or F1 over dimensions (Favreau et al., 2022).
Ranking and retrieval metrics for deployment.
For systems that must order arguments, pairwise and listwise metrics dominate. Pairwise Accuracy (PAcc) on ConvArg and IBM-Pairs remains the primary measure of relative convincingness (Habernal and Gurevych, 2016). Retrieval-oriented work adds MRR and NDCG@$k$; PageRank-style relevance improved NDCG by 4–5 points and is evaluated with Spearman $\rho$ against human rankings (Wachsmuth et al., 2017b). Efficient pairwise annotation plus ranking loss further increases Kendall $\tau$ while halving labelling cost (Gienapp et al., 2020).
System-level and emerging metrics.
Recent studies measure AQA inside larger pipelines. The COVERAGE metric scores how thoroughly a summary reflects argumentative key points without gold labels, serving as an unsupervised proxy for sufficiency (Khosravani et al., 2024). In the AMELIA benchmark for Italian law, system impact is quantified by the rise in judge agreement when automatically ranked arguments are added to briefs (Grundler et al., 2024). Soft-label evaluation—comparing predicted distributions to annotator distributions—operationalises Plank’s call to treat disagreement as signal (Plank, 2022).
5.4 Applications and Future Works
AQA has moved beyond proof-of-concept studies and now underpins a growing range of downstream applications.
Quality-aware retrieval and recommendation.
Search engines that index argumentative content (e.g. args.me) already re-rank results by predicted cogency, sufficiency or style (Wachsmuth et al., 2017b). IBM Debater uses pairwise convincingness scores to surface the strongest counter-points in debate preparation (Gretz et al., 2020). Empirical studies show that integrating AQA scores into BM25 or dense-retrieval pipelines raises NDCG@$5$ without additional user feedback (Gienapp et al., 2020).
Writing-support and educational feedback.
Automated writing-evaluation tools such as ArgRewrite provide sentence-level suggestions that target logical sufficiency and rhetorical clarity; classroom deployments report statistically significant gains in argumentative essay scores (Favreau et al., 2022). In science-education contexts, mapping Walton’s Critical Questions to teacher feedback helps novices revise reasoning chains more effectively (Winkle et al., 2020).
Professional and domain-specific use cases.
The legal benchmark AMELIA attaches sufficiency and statute-relevance scores to Italian court opinions, enabling intelligent brief-generation and judge-support systems (Grundler et al., 2024). In corporate governance, argument-mapping dashboards highlight reasoning gaps in environmental, social and governance (ESG) reports, assisting compliance officers (Palmieri, 2024).
Research Gap and Future works
LLM-based judging offers state-of-the-art accuracy, free natural-language justifications, and sensitivity to emotional bias, but it also inherits dataset biases and incurs computational cost. Future work must explore adversarial test suites, distributional labels, and cost-aware cascading architectures to harness LLM strengths without sacrificing robustness.
6 Modelling Paradigms in the LLM Era
Large Language Models (LLMs) have transformed argument mining from a pipeline of narrow, task-specific classifiers into a set of interaction paradigms in which one or more foundation models collaborate with retrieval modules, symbolic structures or lightweight adapters. Rather than ask whether an architecture is “better” than another, current research asks how LLMs are enlisted as zero-shot predictors, few-shot imitators, knowledge retrievers, judges, data generators, or all of the above. This section synthesises the dominant paradigms, emphasising the design rationales, empirical pay-offs, and unresolved pitfalls.
6.1 Prompt-based Argument Mining
Early demonstrations showed that even naïve declarative prompts (“Is the following sentence a claim?”) enable GPT-3 to match feature-rich baselines on component segmentation and stance detection without gradient updates (Chen et al., 2024). Subsequent work moved beyond template engineering to embed task structure directly inside the prompt. Altemeyer et al. (2025) inserted chain-of-thought (CoT) exemplars that model the logical leap between premise and conclusion; the reasoning trace, rather than the final label, accounted for most of the performance gain on sufficiency and convincingness. Prompting can also express task hierarchy: the PITA framework conditions relation prediction on previously elicited component spans within the same conversation context and reports significant improvements on Reddit and legal briefs (Sun et al., 2024b). Role prompts further reduce topical and demographic bias: framing the model as a Debate Judge or Devil’s Advocate attenuates polarity skew in political fact-checking (Dayanik and Padó, 2020).
Yet prompt brittleness persists. Small lexical tweaks can flip stance labels, and CoT traces lengthen latency in interactive writing tools. Robust prompt ensembles and self-consistency voting help but introduce computational overhead; empirical studies suggest latency grows almost linearly with ensemble size (Zhang et al., 2025). Designing prompts that capture argumentation theory while remaining compact therefore remains an open problem.
6.2 Parameter-efficient Instruction and Multi-task Tuning
Zero-shot prompting rarely suffices for domain-sensitive genres such as biomedical abstracts or statutory law. Instruction tuning—fine-tuning on a mixture of task-specific exemplars framed as natural-language instructions—bridges much of this gap. A single T5-base model (Chung et al., 2024) jointly instruct-tuned on claim segmentation, relation prediction and quality scoring yields consistent gains across all three tasks on IAM and ARIES, while adapters keep GPU memory within budget (Cheng et al., 2022; Gemechu et al., 2024). In the legal domain, ArgLegalSumm couples instruction tuning with an argument salience re-ranker, improving summary coherence in Italian case law (Elaraby et al., 2023). Cross-domain continual learning is more fragile: when a biomedical corpus is appended late in the schedule, earlier gains on social-media stance detection erode, echoing catastrophic forgetting observed in multi-task translation (Pang et al., 2024).
Recent work therefore explores parameter-efficient updates: LoRA (Hu et al., 2022), prefix-tuning and bit-fit adapters enable domain adaptation with two orders of magnitude fewer trainable parameters while preserving general AM capability (Li et al., 2024b). Still unsettled is how to balance tasks of differing label density: curriculum strategies that start with coarse component labels before introducing noisy quality scores appear promising, but systematic ablation is scarce.
6.3 Retrieval-Augmented Generation and Reasoning
LLMs remain constrained by context windows and training cut-off dates. Retrieval-augmented generation (RAG) mitigates both limits by injecting external evidence into the prompt. In debate summarisation, RAG systems retrieve topically relevant counter-arguments and background facts, reducing hallucination and increasing factual density in abstractive summaries (Syed et al., 2023). The OpenDebateEvidence benchmark pairs fifteen million statements with evidence links, enabling end-to-end retrieval-generation pipelines that outperform text-only baselines by a meaningful margin (Roush et al., 2024). Retrieval can also be argument-theoretic: semantic-graph matching over Kialo debate trees yields higher recall of undercutters and mitigations than purely lexical search (Ruiz-Dolz et al., 2023).
Several challenges loom. Evidence ranking often inherits popularity biases, favouring mainstream sources over minority viewpoints. Long evidence snippets may eclipse the claim inside the prompt, forcing heuristic truncation. Finally, retrieval latency threatens interactive applications; hybrid dense-sparse indexing and passage caching alleviate but do not solve the issue.
6.4 Synthetic Supervision and Data Augmentation
LLMs themselves can generate training data, especially where expert annotation is scarce. Wagner et al. (2025) show that LLM-synthesised stance corpora, when filtered for diversity, raise low-resource performance more than in-context few-shot learning. Similarly, LLM-generated key-point coverage scores serve as weak labels for automatic evaluation metrics, correlating well with human judgements (Khosravani et al., 2024). Synthetic supervision also facilitates controlled experiments: topic-balanced sampling can equalise stance priors across political entities, supporting fairness audits (Mehrabi et al., 2022).
The downside is distributional drift. Synthetic arguments tend to be lexically fluent, hedge-free and uniformly polite, properties that can over-regularise fine-tuned models. Provenance cards documenting prompt templates, sampling temperatures, rejection rates and post-editing guidelines are therefore being proposed as a best practice (DBLP:conf/facc/Paullada2021). Legal uncertainty over copyright of machine-generated text remains unresolved.
7 Future Directions
As argument mining transitions into the era of large language models (LLMs), a wide range of opportunities emerge alongside challenges that must be systematically addressed. This chapter synthesizes key limitations identified across prior sections and outlines promising future research directions spanning technical, theoretical, and practical dimensions.
7.1 Interpretability and Controllability
LLM outputs can be impressively fluent yet logically opaque. Black-box convincingness scores or abstractive summaries are difficult to trust in law, medicine or policy making, where explanations must be auditable and contestable. Promising directions include symbolic scaffolding, in which an LLM populates a Toulmin-style or Walton-style template before emitting a verdict (Lee et al., 2024; Borg and Bex, 2024); rationale-aligned distillation, where model predictions are accepted only if an intermediate reasoning trace satisfies argument-quality criteria (Zhao et al., 2024); and rule-augmented decoding, which filters candidate continuations against critical-question constraints during generation (Palmieri, 2024). Empirical studies show that such structured intermediates improve user trust more than generic saliency maps, yet rigorous human-in-the-loop evaluations are still rare.
7.2 Data Quality, Bias and Hallucination
LLM training corpora contain ideological skew, demographic imbalance and source-credibility noise. These artefacts surface in persuasive writing: emotionally charged but weakly supported arguments are often overrated by LLM judges (Chen and Eger, 2025). Counterfactual calibration and actor masking reduce partisan bias in stance detection (Li et al., 2025; Dayanik and Padó, 2020), but systematic audits across domains remain scarce. Hallucination is the sister problem: retrieval-augmented generation mitigates unsupported claims yet inherits popularity and language biases from search engines (Roush et al., 2024). Future benchmarks should combine bias slices e.g. position papers from under-represented groups— with factual stress tests to quantify both issues jointly.
7.3 Generalisation and Robustness
Most LLM studies report single-domain results; cross-genre transfers often collapse. Debate summarisers tuned on English legal cases lose more than a third of their content fidelity when applied to multi-party chat (Zhao et al., 2023). Adversarial paraphrase attacks flip stance predictions with minimal lexical edits (Gienapp et al., 2020). Potential remedies include curriculum-style multi-task fine-tuning, meta-learning over domain clusters, and evaluation suites that mix genres, languages and modalities within the same test fold.
7.4 Synthetic Supervision and Benchmark Integrity
LLM-generated corpora such as ZeroStance (Zhao et al., 2024) and OpenDebateEvidence (Roush et al., 2024) accelerate data growth but raise three red flags: (i) contamination models may re-encounter their own pre-training shards; (ii) superficial reasoning synthetic counter-claims are stylistically rich yet logically shallow; (iii) blind spot edge cases rarely appear in temperature-controlled sampling. Hybrid pipelines that blend human editing, diversity-aware sampling and automatic quality-metrics (e.g. coverage and exhaustiveness scores from Khosravani et al. (2024)) are a promising compromise. Benchmark leaders should publish provenance cards documenting prompt templates, sampling temperatures and rejection rates.
7.5 Theory-Grounded Modelling
Most LLM pipelines treat argument quality as a monolithic scalar. Argumentation theory, however, differentiates logical sufficiency, dialectical relevance and rhetorical effectiveness. Embedding these distinctions into prompts, loss functions or evaluation metrics is still nascent. Early attempts to weave Walton’s critical questions into dialogue evaluation improved error localisation but slowed inference (Palmieri, 2024). Future work could explore neuro-symbolic hybrids that use graph-based logic checkers as fast filters before LLM re-drafting.
7.6 Human-Centred and Interactive AM
Educational feedback tools, civic debate platforms and policy consultation interfaces require models that operate with users, not merely for them. That includes tracking my-side bias (Baccini and Hartmann, 2022), adapting critique to a learner’s knowledge state (Diana et al., 2020), and providing actionable, value-aligned suggestions. Reinforcement learning from argumentative feedback, real-time uncertainty estimates and personalisation via retrieval of prior user arguments are promising building blocks.
8 Conclusion
The field of Argument Mining has come a long way, moving from narrow, feature-engineered pipelines trained on small, task-specific corpora to a rich ecosystem powered by large language models. Today’s systems can identify claims, map support and attack relations, gauge stance, assess quality, and even summarize key points, all within a unified framework that leverages prompting, few-shot adaptation, retrieval of background evidence, and synthetic data generation. Alongside these gains, however, have come new challenges: the opacity of model reasoning, the risk of bias and hallucination, brittleness under domain shift, and questions of provenance and ethical deployment.
Looking ahead, progress will hinge on our ability to reconcile raw performance with transparency, robustness and theoretical grounding. This means developing interpretable abstractions e.g. symbolic scaffolds, rationale-based decoding and uncertainty-aware metrics, that expose the “why” behind LLM decisions; constructing datasets and benchmarks that capture the full spectrum of argument modalities, languages and genres; auditing for demographic, ideological and stylistic bias; and designing interactive, human-centric systems that guide rather than replace expert judgment. By weaving together advances in model technology, data design and evaluation practice, we can build argument-mining tools that are not only accurate, but also trustworthy, equitable and broadly useful in the real world.